# : Towards Semantic Steganography via Large Language Models

### v1

@@ Abstract
Although steganography has made significant advancements in recent years, it still struggles to embed semantically rich, sentence-level information into carriers. However, in the era of AIGC, the capacity of steganography is more critical than ever. In this work, we present Sentence-to-Image Steganography, an instance of Semantic Steganography, a novel task that enables the hiding of arbitrary sentence-level messages within a cover image. Furthermore, we establish a benchmark named Invisible Text (IVT), comprising a diverse set of sentence-level texts as secret messages for evaluation. Finally, we present : Semantic Steganographic Language Model, which utilizes large language models (LLMs) to embed high-level textual informationβsuch as sentences or even paragraphsβinto images. Unlike traditional bit-level counterparts, enables the integration of semantically rich content through a newly designed pipeline in which the LLM is involved throughout the entire process. Both quantitative and qualitative experiments demonstrate that our method effectively unlocks new semantic steganographic capabilities for LLMs. The source code will be released soon.
@@ 1 Introduction
In an age long past, under the shadow of the Persian Empire, Histiaeus, ruler of Miletus, devised a plan that would etch his name in the annals of history. Under the watch of the imperial, he tattooed a secret message onto the scalp of a loyal slave, waiting for the hair to regrow and conceal the message beneath a veil of innocence. When the time was right, the slave crossed enemy lines, carrying the message not in his hands but under his skin. The message sparked a revolution, a call to arms that would ignite the Ionian Revolt. This tale, immortalized by Herodotus in the Histories, remains one of the earliest recorded feats of Steganography - the art of hiding information in plain sight.
Fast forward today, on the verge of entering the Artificial Intelligence Generated Content (AIGC) era, the principles of steganography have never been more relevant, with the prosperity of various generative models (e.g., Stable Diffusion [rombach2021highresolution], ChatGPT [openai_chatgpt], and Llama [touvron2023llama]), which can produce highly realistic text [bai2023qwentechnicalreport, qwen2025qwen25technicalreport], images [rombach2021highresolution, ramesh2022hierarchicaltextconditionalimagegeneration] and videos [liu2024sorareviewbackgroundtechnology], and even execute tasks with human-like proficiency [durante2024agentaisurveyinghorizons]. Their widespread application has blurred the line between fact and forgery, raising numerous security and legal concerns. The application of the generative model may foster the generation of biased content [ling2025bias, serouis2024exploring, kumar2025no], image manipulation [qu2024towards], and the creation of fake news [wu2024fake]. Moreover, the human-like content generated by those models is flooding the entire Internet, seriously hindering the determination of the credibility of evidence in forensics [yin2025digital]. Therefore, there is an urgent need for effective measures to prevent the large-scale misuse of generative models.
To address this problem, embedding forensics information into the generated content by steganography could be the most effective method. As the real-world application of steganography, embeds secret information that is typically difficult to detect and remove but can be easily retrieved [hartung2002multimedia]. With this technique, it is possible to determine whether the generative model creates the content quickly. Moreover, the extracted information can also be used for copyright identification and source tracing, thereby protecting intellectual property rights and helping locate the infringer. Nowadays, various strategies have been developed for text [pan2024markllm, zhang2024remark, liang2024watermarkingtechniqueslargelanguage], images [wen2023tree, fernandez2023stable], video [lin2024automatic, luo2023dvmark], and audio [zhang2024v2a] generated by generative models.
As a primary medium of visual communication, images are widely shared on the Internet, promoting the development of digital image steganography. Over years of research, numerous methods have been proposed[zhu2018hidden, wengrowski2019light, tancik2020stegastamp, wen2023tree, cox2007digital, wang2001image, patel2012steganography]. However, most existing approaches focus on embedding low-level binary or character-based information, neglecting the semantic relationships inherent in natural language. Consequently, these methods exhibit limited capacity and poor expressivity, preventing them from embedding sentence-level messages. Those methods underutilize the representational potential of modern high-resolution images and fail to meet the increasing capacity demands of watermarking in the era of generative AI.
@@
For further analysis, we divide the current steganography [zhu2018hidden, wengrowski2019light, tancik2020stegastamp, patel2012steganography] pipeline into three processes: the encoding process, which transforms the secret information into a suitable representation; the insertion process, which embeds the message into the cover image to produce a stego image; and the decoding process, which extracts the secret message from the stego image. As illustrated in Figure 1, three significant challenges hinder current steganographic methods from embedding substantial semantic information into images: 1) In the encoding process, secret messages are typically converted into bitstreams based on ASCII or UTF formats. This method is simple but primitive, neglects the semantic information between words, has a low information density, and is unsuitable for encoding large information such as sentences; 2) In the insertion process, the encoder takes the secret message and cover image as input to generate a stego image. The interaction between two fundamentally different modalities (bitstreams and images) results in suboptimal insertion performance. 3) In the decoding process, due to the aforementioned limitations, the decoder can only recover bitstreams, restricting the hidden information to the word level. The three significant challenges mentioned above hinder the existing steganography to increase the secret message capacity to include semantic information, such as complete sentences.
@@
To address these challenges, we introduce the broader concept of Semantic Steganography, which focuses on embedding semantically meaningful and structured contentβsuch as sentences or paragraphsβinto carriers. Building on this idea, we formalize the Sentence-to-Image Steganography task as an instance of semantic steganography. This task requires hiding sentence-level or even paragraph-level semantic information into a cover image. Notably, the secret message can consist of arbitrary textual content of any length. To accomplish this task, the model must possess two key capabilities: 1) encoding large and complex information into an appropriate representation and 2) effectively handling two fundamentally different modalities (text and image).
Recent advancements in the development of large language models (LLMs) have spurred efforts to integrate LLMs with vision backbones, leading to the rise of vision-language models (VLMs) [zhang2024vision, ye2024mplug, li2022mplug, liu2024improved, liu2023visual]. However, most of these models focus on vision understanding and text generation. This raises an important question: Can LLMs be used directly as a universal steganographic tool to embed sentence-level information into carriers such as images?
To this end, we propose the Semantic Steganographic Language Model (), an innovative steganographic framework. To the best of our knowledge, this is the first work to explore hiding sentence-level and even paragraph-level information in images. Specifically, we redesign the entire pipeline used in current watermarking strategies. As illustrated in FigureΒ 1, we first encode the secret sentence into tokens using a tokenizer. Second, during the insertion process, we use the LLM to generate secret embeddings without relying on image priors, thereby eliminating interference between the cover image and the secret message. Finally, we use the same LLM to decode the secret message from the stego image.
In our empirical experiments, can conceal up to 500 words of text into a single image with a resolution of , breaking the boundary between image and text. Additionally, we evaluate the steganographic performance of different LLMs within the framework and verify its versatility across various language models.
@@
We propose a new steganographic taskβSentence-to-Image Steganography, which requires embedding sentence-level information into a cover image.
@@
We establish a benchmark for the semantic steganography named Invisible Text (IVT), containing thousands of sentence-level samples from diverse sources. This benchmark is essential for evaluation and encourages the community to further explore the potential of semantic steganography.
2 Related Works
@@ 2.1 Image Steganography
Steganography is a technique for concealing secret information within a carrier while preserving its visual appearance and avoiding detection. Traditional image steganography methods, such as the Least Significant Bit (LSB) algorithm [wang2001image] and frequency-domain approaches [patel2012steganography, chen2006dwt] (e.g., DCT and DWT), embed secret data by modifying the pixel values or transform coefficients of the cover image. However, these methods have limited capacity and struggle to meet the requirements of modern steganography, whereas deep learning-based methods perform better. Recent advances in deep learning have produced more powerful steganographic techniques, such as HiDDeN [zhu2018hidden], LFM [wengrowski2019light], and StegaStamp [tancik2020stegastamp], which can embed bit-level information in images with robustness against various manipulations. Moreover, methods like HiNet [jing2021hinet], ISN [lu2021large], LanNet [lan2023robust], and StegFormer [ke2024stegformer] offer greater capacity, enabling the concealment of an image in carrier with the same resolution.
@@ 2.2 Large Language Models in Computer Vision
The prosperity of the current large language model (LLM) has pushed the whole community of the Natural Language Processing (NLP) into new boundaries. The computer vision community has also noticed the tremendous success of the practice of LLM. Consequently, lots of the latest research focuses on exploring the integrity of the LLM with the vision backbone. BLIP [li2022blip] proposes a Multi-modal mixture of Encoder-Decoder (MED), while BLIP2 [li2023blip] introduces a frozen LLM to bootstrap vision-to-language learning, and proposes a Q-Former to bridge the modality gap. LLaVA [liu2023visual] connects the vision backbone and Llama [touvron2023llama] and tunes the LLM with multi-modal language-image instruction-following data, achieving the general-purpose visual and language understanding.
The methods above explore the LLM application in visual understanding, often seen as a high-level vision task. However, the LLM can also enhance the performance of the low-level vision backbone. LISA [Lai2023LISARS] uses a novel segmentation token and unlocks the segmentation capability of the LLM, which can be used for reasoning segmentation tasks. Xu et al. used localized image captioning data to pretrain PixelLLM to equip it with localization capability [xu2024pixel]. Zheng et al. froze LLM to solve a range of low-level vision tasks, demonstrating the LLMβs strong generalization capability [zheng2024lm4lv].
In this paper, we investigate the steganographic capabilities of LLMs, which has not been explored in the previous works. Experiments show that LLMs can serve as a steganographic tool to embed sentence-level information into a cover image.
3 Sentence-to-Image Steganography
The whole steganographic community has traditionally focused on low-level information hiding (e.g., at the bit or image level), overlooking the semantic structure of the hidden information. In contrast, we introduce a new direction of Semantic Steganography, which refers to hiding semantically meaningful and structured content, such as sentences or paragraphs, in the carrier. This paradigm redefines what kind of information can be hidden. Rather than merely embedding arbitrary bitstreams, it enables directly hiding and recovery of natural language messages, which benefits the steganographic techniques by expanding their expressive and potential use cases. Specifically, we formalize the Sentence-to-Image Steganography task, which aims to embed sentence-level (or even paragraph-level) natural language message into images such that the stego image remains visually indistinguishable from a benign image while allowing faithful recovery of the original message.
@@ 3.1 Problem Definition
The sentence-to-image steganography task aims to embed a secret message into a cover image to produce a stego image , from which can be fully recovered. This task emphasizes high capacity, seeking to conceal as much meaningful message as possible without degrading the visual quality of the cover image. Moreover, the secret messages can be any sentence-level, even paragraph-level data, encompassing more intricate expressions and longer text.
@@ 3.2 Benchmark
However, there is an absence of datasets specially designed for evaluating semantic steganography. To fill this gap, we construct the Invisible Text (IVT) dataset, which comprises a diverse set of sentence-level texts drawn from existing text datasets and generated by LLMs, serving as secret messages. We divide the messages into three levels of granularity: short messages (one sentence), medium messages (two to three sentences), and long messages (full articles), denoted as IVT-S, IVT-M, and IVT-L, respectively. Please refer to Appendix A for more details about the IVT dataset.
4 Our Method
4.1 Representation Information as Tokens
The representation of secret text is a critical aspect of semantic steganography. Previous methods convert each character of a word into bits using UTF encoding[tancik2020stegastamp, zhu2018hidden, wengrowski2019light]. However, these primitive approaches have two significant limitations. First, UTF encoding lacks semantic information, which is essential for representing sentence-level content. Second, representing a sentence as a bitstream is inefficient, as the resulting sequence becomes excessively long and difficult to process. We hypothesize that such primitive representations are particularly fragile when encoding richer semantic content. Therefore, it is necessary to explore more effective representations for sentence-level information.
To address these issues, we leverage the tokenizer of the LLM to convert the secret sentence into tokens. Tokenization is a widely used technique in NLP, where sentences are broken down into smaller units called Tokens. Tokens capture richer semantic information and facilitate modeling relationships between words. Moreover, tokenization significantly reduces sequence length, which benefits downstream processing. Using the tokenizer, we represent the secret sentence as a sequence of tokens. This approach not only preserves the semantic structure of the sentence but also ensures an efficient representation for embedding.
@@ 4.2 LLMs for Steganography
As shown in FigureΒ 2 the consists of three key modules: a pre-trained LLM for embedding and decoding the secret information, a Token-to-Patch MLP for aligning text features and image feature, and a Patch-to-Token MLP for embedding visual features to token embeddings. Notably, is a universal framework in which the LLM can be ANY pre-trained LLM.
Semantic embedding via LLMs. The LLM is recognized as the de facto standard architecture in the NLP community. Recent studies, however, have demonstrated that LLM can serve as a universal interface with different instructions in language to complete different tasks. In the computer vision community, researchers have explored integrating LLMs with vision backbones for segmentation, localization, and other visual tasks. Inspired by these works, we employ an LLM to generate embeddings conditioned by secret information, which will be inserted into the cover image to produce the stego image. Owing to the strong generalization ability of LLMs, we are able to compress the semantic information into a few embeddings.
Specifically, as illustrated in Figure 2, we design an embedding prompt and concatenate it with a secret message to instruct the LLM to embed the secret message into embedding, which we call secret message embeddings (SMEs). We call this strategy Prompt-Guide Embedding. The SMEs are then aligned with the image feature using the . After that, we reshape it and add to the cover image to get a stego image.
Universal insertion. To insert SMEs into the cover image, we use to align the token embeddings with the image features. Then, we adapt the UDH method [zhang2020udh], which directly adds the SMEs to the cover image to produce a stego image. Previous methods generate the stego image conditioned on both the cover image and the secret message. Empirically, however, the UDH method decouples the secret message from the carrier, resulting in smoother training and improved performance.
@@
Template designing As illustrated in Figure 3, we carefully design the embedding and decoding prompts to instruct the LLM to embed and recover the secret message. We introduce two special tokens: and , which are used to separate secret messages from the prompt. We use the output embeddings of the LLM as SMEs. For clarity, we denote the location of the SMEs in the output of LLM as . During the decoding process, the extracted feature will be concatenated with the decode prompt as the input of LLM. We use as the placeholder for the feature extracted from stego image by .

### v2

@@ Abstract
Despite remarkable progress in steganography, embedding semantically rich, sentence-level information into carriers remains a challenging problem. In this work, we present a novel concept of Semantic Steganography, which aims to hide semantically meaningful and structured content, such as sentences or paragraphs, in cover media. Based on this concept, we present Sentence-to-Image Steganography as an instance that enables the hiding of arbitrary sentence-level messages within a cover image. To accomplish this feat, we propose : Semantic Steganographic Language Model, which leverages large language models (LLMs) to embed high-level textual information into images. Unlike traditional bit-level approaches, redesigns the entire pipeline, involving the LLM throughout the process to enable the hiding and recovery of arbitrary sentences. Furthermore, we establish a benchmark named Invisible Text (IVT), comprising a diverse set of sentence-level texts as secret messages to evaluate semantic steganography methods. Experimental results demonstrate that effectively enables direct sentence recovery beyond bit-level steganography. The source code and IVT dataset will be released soon.
@@ 1 Introduction
Steganography is the art of hiding information in plain sight. Historically, it was used as a means of covert communication during wartime. Nowadays, these techniques have found broader applications in copyright protection [80], tamper detection [79, 81], and secure communication [72].
With the advent of deep learning, steganography has achieved remarkable progress in recent years. However, despite this apparent prosperity, existing steganographic models can hide only few words of secret information. This limitation primarily arises from their design paradigm: the secret message must first be converted into a bitstream before being processed by the model. Consequently, current methods are fundamentally designed to embed binary data [68, 60, 9], which severely restricts their capacity and leads to low communication efficiency in practical applications.
@@
For further analysis, we take image steganography as a case study and divide the steganographic pipeline into three processes, as illustrated in Fig.Β 1. There are three significant limitations that hinder existing methods from embedding larger amounts of information into images: 1) In the encoding process, messages are converted into bitstreams via ASCII or UTF encoding. This method is simple but primitive, neglects the semantic relationships between words, has a low information density, and is unsuitable for encoding large amounts of information, such as sentences. 2) In the insertion process, the encoder takes the secret message and cover image as input to generate a stego image. The interaction between two fundamentally different modalities (bitstreams and images) results in suboptimal insertion performance. 3) In the decoding process, due to the aforementioned constraints, the decoder can only recover bitstreams, restricting the hidden information to the word level. These practical limitations call for a rethinking of steganography: Are steganographic frameworks inherently limited to hiding low-level bitstream information?
@@
Motivated by this question, we introduce the broader concept of Semantic Steganography, which focuses on embedding semantically meaningful and structured content, such as natural language, into carriers. Building on this idea, we formalize the Sentence-to-Image Steganography task as an instance of semantic steganography. This task requires hiding sentence-level or even paragraph-level information in a cover image. Notably, the secret message can consist of arbitrary textual content of any length. To accomplish this task, the model must possess two key capabilities: encoding large and complex information into an appropriate representation and effectively handling two fundamentally different modalities.
Recent advancements in large language models (LLMs) have spurred research on integrating them with vision backbones, leading to the rise of vision-language models (VLMs) [76, 70, 37, 42, 43]. These studies demonstrate that LLMs exhibit strong capabilities in semantic understanding and cross-modal reasoning, which precisely align with the requirements of semantic steganography. This alignment highlights the potential of LLMs to serve as a foundation for developing a universal steganographic framework that embeds high-level semantics within carriers.
Building on this insight, we propose the Semantic Steganographic Language Model (), a novel framework that leverages the semantic understanding capability of LLMs for steganography. To address the three challenges mentioned above, completely redesigns the entire pipeline of current methods, as illustrated in Fig.Β 1. 1) In the encoding process, we first transform the secret textual information into a sequence of discrete tokens using a tokenizer. This process not only captures the semantic content of the sentence but also preserves the relationships between words, enabling more effective and structured representation of the secret message. 2) During the insertion process, we leverage LLM to generate secret embeddings without relying on any prior of the cover image. By decoupling secret embeddings from the carrier, prevents potential interference between the hidden message and the cover image. 3) In the decoding process, we employ the same LLM to decode the secret message from the stego image. Overall, provides a unified framework to handle the sentence-to-image task, demonstrating strong capability in encoding and recovering semantically rich messages without compromising visual integrity. In our qualitative experiments, successfully conceals up to 500 words of text into a single image with a resolution of , breaking the boundary between image and text. We further evaluate the steganographic performance of different LLMs within the framework and verify its versatility across various models.
@@
We propose the concept of Semantic Steganography and formalize Sentence-to-Image Steganography as an instance, redefining the traditional goal of steganography from embedding low-level data to concealing meaningful language content within cover media.
@@
We establish a benchmark for semantic steganography named Invisible Text (IVT), containing thousands of samples from diverse sources. This benchmark is essential for evaluation and encourages the community to further explore the potential of semantic steganography.
2 Related Work
@@ 2.1 Image Steganography
Steganography is a technique for concealing secret information within a carrier while preserving its visual appearance and avoiding detection. Traditional image steganography methods, such as the Least Significant Bit (LSB) algorithm [65] and frequency-domain approaches [48, 12], embed secret data by modifying the pixel values or transform coefficients of the cover image. However, these hand-crafted methods tend to introduce statistical artifacts that can be easily detected by steganalysis, whereas deep learning-based methods achieve higher security. Recent advances in deep learning have produced more powerful steganographic techniques, such as HiDDeN [83], LFM [68], and StegaStamp [60], which can embed bit-level information in images with robustness against various manipulations. Moreover, methods like HiNet [29], ISN [45], LanNet [36], and StegFormer [31] offer greater capacity, enabling the concealment of an image in a carrier with the same resolution. Recently, several works have explored utilizing diffusion models for steganography, such as DGADM-GIS [73], RoSteALS [9], and Latent Watermark [46].
@@ 2.2 Large Language Models in Computer Vision
The prosperity of current LLMs has pushed the entire Natural Language Processing community into new boundaries. The computer vision community has also recognized the tremendous success of LLMs. Consequently, lots of the latest research focuses on exploring the integrity of the LLM with the vision backbone. BLIP [39] proposes a Multi-modal mixture of Encoder-Decoder (MED), while BLIP2 [38] introduces a frozen LLM to bootstrap vision-to-language learning, and proposes a Q-Former to bridge the modality gap. LLaVA [43] connects the vision backbone and Llama [62] and tunes the LLM with multi-modal language-image instruction-following data, achieving the general-purpose visual and language understanding.
The methods above explore the LLM application in visual understanding, often seen as a high-level vision task. However, the LLM can also enhance the performance of the low-level vision backbone. LISA [35] uses a novel segmentation token and unlocks the segmentation capability of the LLM, which can be used for reasoning segmentation tasks. Xu et al. used localized image captioning data to pretrain PixelLLM to equip it with localization capability [69]. Zheng et al. froze LLM to solve a range of low-level vision tasks, demonstrating the LLMβs strong generalization capability [82].
While several studies have attempted to apply LLMs to linguistic steganography [66, 5, 54], most of them rely on manually designed sampling strategies. In this paper, we integrate LLM into the entire steganographic pipeline, directly utilizing LLM to generate secret embeddings directly, and successfully scale the steganographic capacity from bit to sentence-level.
3 Semantic Steganography
The whole steganographic community has traditionally focused on low-level information hiding (e.g., at the bit or image level), overlooking the semantic structure of the hidden information. In contrast, we introduce a new direction of Semantic Steganography, which refers to hiding semantically meaningful and structured content, such as sentences or paragraphs, in the carrier. This paradigm redefines what kind of information can be hidden. Rather than merely embedding arbitrary bitstreams, it enables the direct hiding and recovery of natural language messages, which benefits the steganographic techniques by expanding their expressive and potential use cases.
@@ 3.1 Problem Definition
We define semantic steganography as the task of hiding semantic content within a carrier while keeping it undetectable. In this work, the semantic content of interest is expressed in natural language, specifically as complete sentences that convey coherent and interpretable meaning.
Formally, let denote a natural language message, and denote a cover object. The encoder embeds into , producing the stego object . The corresponding decoder aims to recover the original message as . The semantic steganography task must satisfy the following constraints:
Specifically, recoverability measures the semantic consistency of the recovery message using the similarity function . Imperceptibility evaluates the similarity between cover and stego objects, where measures perceptual distortion, and denotes the maximum allowable perceptual difference. Security reflects the indistinguishability between cover and stego objects against steganalysis. represents the output of the steganalysis tool.
In essence, semantic steganography generalizes the notion of information hiding from the signal level to the semantic level. By encoding linguistic meaning rather than raw bits, it establishes a unified framework for embedding natural language semantics across arbitrary modalities.
@@ 3.2 Benchmark
However, there is an absence of datasets specially designed for evaluating semantic steganography. To fill this gap, we construct the Invisible Text (IVT) dataset, which comprises a diverse set of sentence-level texts drawn from existing text datasets and generated by LLMs, serving as secret messages. We divide the messages into three levels of granularity: short messages (one sentence), medium messages (two to three sentences), and long messages (full articles), denoted as IVT-S, IVT-M, and IVT-L, respectively. Please refer to Sec.Β 7 of supplementary materials for more details.
3.3 Sentence-to-Image Steganography
Specifically, we formalize the sentence-to-image steganography task as a concrete instantiation of the semantic steganography. Following the definition above, the object in this task is specified as an image . The objective is to secretly embed natural-language messages within the image while ensuring that the hidden information can be accurately recovered.
4 Method
4.1 Representing Information as Tokens
The representation of secret message is a critical aspect of semantic steganography. Previous methods convert each character of a word into bits using UTF encoding[60, 83, 68]. However, such character-level representations have two significant limitations. First, UTF encoding lacks semantic compression capability, treating each character equally regardless of its meaning. Second, encoding a sentence as a bitstream leads to an excessively long sequence that is difficult to process. We hypothesize that such primitive representations are fragile when encoding richer semantic content. Therefore, it is necessary to explore more effective representations for sentence-level information.
To address these issues, we leverage the tokenizer to convert the secret sentence into tokens. Tokenization [67] is a widely used technique in NLP, where sentences are broken down into smaller units called Tokens. Unlike character-level encoding, tokens correspond to semantically meaningful units, such as words, sub-words, or phrases, allowing a more compact representation. This not only preserves the semantic structure of the sentence but also significantly shortens the sequence length, improving the efficiency and stability of downstream embedding and decoding processes.
@@ 4.2 LLMs for Steganography
As shown in Fig.Β 2, the consists of three key modules: a pre-trained LLM for embedding and decoding the secret information, a Token-to-Patch MLP for aligning text features and image features, and a Patch-to-Token MLP for embedding visual features to token embeddings. Notably, is a universal framework in which the LLM can be ANY pre-trained LLM.
Semantic Embedding via LLMs. The LLM is recognized as the de facto standard architecture in the NLP community. Recent studies, however, have demonstrated that LLM can serve as a universal interface with different instructions in language to complete different tasks. In the computer vision community, researchers have explored integrating LLMs with vision backbones for segmentation [35], grounding [69, 50, 30], and other visual tasks [74]. Inspired by these works, we employ an LLM to generate embeddings conditioned by secret information, which will be inserted into the cover image to produce the stego image. Owing to the strong generalization ability of LLMs, we are able to compress the semantic information into a few embeddings.
Specifically, as illustrated in Fig.Β 2, we design an embedding prompt and concatenate it with a secret message to instruct the LLM to embed the secret message into embeddings, which we call secret message embeddings (SMEs). We call this strategy Prompt-Guide Embedding. The SMEs are then aligned with the image feature using the . After that, we reshape it and add to the cover image to get a stego image.
Universal Insertion. To insert SMEs into the cover image, we use to align the token embeddings with the image features. Then, we adapt the UDH method [75], which directly adds the SMEs to the cover image to produce a stego image. Previous methods generate the stego image conditioned on both the cover image and the secret message. Empirically, however, the UDH method decouples the secret message from the carrier, resulting in smoother training and improved performance.
@@
Template Designing. As illustrated in Fig.Β 3, we carefully design the embedding and decoding prompts to instruct the LLM to embed and recover the secret message. We introduce two special tokens: and , which are used to separate secret messages from the prompt. We use the output embeddings of the LLM as SMEs. For clarity, we denote the location of the SMEs in the output of LLM as . During the decoding process, the extracted feature will be concatenated with the decode prompt as the input of LLM. We use as the placeholder for the feature extracted from stego image by .